I need help on setting this problem up!? "A rectangular garden is fenced on 3 sides with a wall forming the 4th side. The total length of the fence is 30m. The area of the garden is 112m squared. Find the dimensions of the garden."
Answer by ankor@dixie-net.com(22740)

A rectangular garden is fenced on 3 sides with a wall forming the 4th side.
The total length of the fence is 30m.
The area of the garden is 112m squared.
Find the dimensions of the garden."
:
Since the wall forms one side the fence equation
L + 2W = 30
L = (30-2W); use this form for substitution
:
The Area
L * W = 112
Replace L with (30-2W)
(30-2W) * W = 112
We can form a quadratic equation from this
-2W^2 + 30w - 112 = 0
simplify and change the sign, divide by -2
W^2 - 15w + 56 = 0
Factors to
(W-8)(W-7) = 0
W = 8 m
W = 7 m
Either solution will work, I'll let you find L for each solution